OpenStudy (anonymous):

Please help I've been trying to figure it out but I just don't get it. WILL MEDAL! Which ordered pair is a solution of y = x - 4? (-1, -5) (0, -3) (2, 6) (3, 1)

OpenStudy (anonymous):

Remember this: (x;y) ! So if (-1,-5) ====> y=-5, x=-1 ===> -5=-1-4 Done :) same with others

OpenStudy (anonymous):

@owllover123

OpenStudy (anonymous):

Hmm

OpenStudy (anonymous):

@iGreen @perl @gottennis121 can any of you help?

OpenStudy (gottennis121):

Ok, first the numbers in the options are x first, then y. Stick in the numbers in the equation, and see if the answer reflects the y. I'll help with the first one. (y) -5 = (x) -1 - 4. So, 4 - -1 = ?

OpenStudy (gottennis121):

Ok, I'll give the answer to the first equation. 4 - -1 = 5. That isn't the answer because you need -5

OpenStudy (anonymous):

umm sorry still confused and how does 4 - -1 = 5?

OpenStudy (gottennis121):

4 is the larger number, so the answer will be positive. then you switch the sign around since this is the case and take out the - from the 1. So, really the equation is 4 + 1

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

then what do you do?

OpenStudy (anonymous):

@MSxSammiBooxX can you help

OpenStudy (gottennis121):

4 - 0 = ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

???

OpenStudy (gottennis121):

So that isn't the answer either because it says that answer (that was 4) should be -3. Let's try the next equation: 4 - 2 = ? Not 6.

OpenStudy (anonymous):

2

OpenStudy (gottennis121):

Try the final equation 4 - 3 = ?

OpenStudy (anonymous):

1

OpenStudy (gottennis121):

There's your answer! 4 - 3 = 1. So what's the answer? 1 (y) = 3 (x) - 4. All I did was switch the equation around

OpenStudy (anonymous):

Is the answer the third one? (2, 6)?

OpenStudy (gottennis121):

No

OpenStudy (anonymous):

(-1, -5)?

OpenStudy (gottennis121):

You told me that 4 - 3 (x) = 1 (y).

OpenStudy (gottennis121):

No. 3 is x and 1 is y. I put the numbers into the equation and it fit. It's D. Do you see why? or no?

OpenStudy (anonymous):

ohhhhhhh okk I get it now! Thanks SO MUCH!

OpenStudy (gottennis121):

no prob

OpenStudy (anonymous):

AHHHH! I just turned in my test it was actually A

OpenStudy (anonymous):

The answer was A

OpenStudy (anonymous):

;-( I got 50%

OpenStudy (gottennis121):

it can't be

OpenStudy (anonymous):

That's what it was